OpenStudy (anonymous):

Decide which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring. −b b2 − 4ac 2a Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 2x2 + 7x + 3 = 0

OpenStudy (anonymous):

if \(b^2-4ac\) is a perfect square, then you can factor using rational numbers

OpenStudy (anonymous):

tell your teacher you can always factor, just not always using integers

OpenStudy (anonymous):

\[2x^2 + 7x + 3 = 0\]\[(x+3)(2x+1)=0\]

OpenStudy (anonymous):

its on a quiz. i have to answer with a numerical answer

OpenStudy (anonymous):

can you solve \((x+3)(2x+1)=0\) ?

OpenStudy (anonymous):

set each factor equal to zero and solve

OpenStudy (anonymous):

for example \[2x+1=0\] \[2x=-1\] \[x=-\frac{1}{2}\]